### v2

# Multilayer network science: theory, methods, and applications

Abstract
Multilayer network science has emerged as a central framework for analysing interconnected and interdependent complex systems. Its relevance has grown substantially with the increasing availability of rich, heterogeneous data, which makes it possible to uncover and exploit the inherently multilayered organisation of many real-world networks. In this review, we summarise recent developments in the field. On the theoretical and methodological front, we outline core concepts and survey advances in community detection, dynamical processes, temporal networks, higher-order interactions, and machine-learning-based approaches. On the application side, we discuss progress across diverse domains, including interdependent infrastructures, spreading dynamics, computational social science, economic and financial systems, ecological and climate networks, science-of-science studies, network medicine, and network neuroscience. We conclude with a forward-looking perspective, emphasizing the need for standardised datasets and software, deeper integration of temporal and higher-order structures, and a transition toward genuinely predictive models of complex systems.
I Introduction
During the last 25 years, network science has established itself as a powerful and seamlessly integrated research discipline Dorogovtsev and Mendes (2013); Vespignani and Pastor-Satorras (2004); Caldarelli (2007); Barrat et al. (2008); Cohen and Havlin (2010); Newman (2018); Estrada (2011); Estrada and Knight (2015); Barabási (2016); Latora et al. (2017); Menczer et al. (2020); Bagrow and Ahn (2024). At variance with other, more niched disciplines, it has allowed us to enrich and to synergize with other branches of science: from sociology and economics to biology, chemistry, and physics. This has been possible because of the data revolution and increased computational power that took place in the twilight of the 20th century. Nowadays, applications of network science tools pervade all fields of science but, most importantly, the discipline has changed the way we think about interactions in natural and human-made complex systems.
The momentum for the development of network science came with the discovery that seemingly diverse networks have universal properties that pervade many natural and artificial systems. Since 1998, ignited by the works of Watts and Strogatz and Barabási and Albert Watts and Strogatz (1998); Barabási and Albert (1999), systems as diverse as electric power grids, food chains, brain networks, protein networks, transcriptional networks, and social networks have been cast into the mathematical language of networks. We are now familiar with jargon that includes centrality, homogeneous and heterogeneous systems, communities, structural and dynamic phase transitions, weak and strong links, etc.
Yet, as the field matured, it became increasingly clear that most real systems cannot be faithfully represented as individual networks. Many natural, social, and technological systems are instead composed of several types of interactions, giving rise to multilayer or interdependent structures Kivelä et al. (2014); Boccaletti et al. (2014); Battiston et al. (2014); Bianconi (2018); Aleta and Moreno (2019); Artime et al. (2022); De Domenico (2023). The conceptual roots of this idea trace back to the sociological notion of multiplexity, where individuals were known to maintain several types of social ties simultaneously within the same population Wasserman and Faust (1994). These early multi-relational studies anticipated many of the ideas later formalized in network science. Then, with the rise of large-scale digital data around 2010, empirical analyses began to demonstrate that distinct layers of interaction can exhibit markedly different structural and dynamical patterns, underscoring the need for explicit multilayer representations Szell et al. (2010).
A major turning point came with the study of interdependent networks, where nodes in one layer depend on nodes in another layer for functionality Buldyrev et al. (2010); Vespignani (2010). This framework revealed that coupling multiple networks can fundamentally alter systemic behaviour, producing abrupt percolation transitions and cascading failures, phenomena not observed in isolated networks. Following these foundational insights, the field rapidly matured from focusing primarily on cascades and robustness to developing a comprehensive mathematical formalism and exploring a wide array of dynamic processes. Researchers worked to create a unified language and a robust mathematical framework to describe the diverse architectures of multilayer systems Kivelä et al. (2014); Boccaletti et al. (2014); Battiston et al. (2014), enabling the extension of classical concepts of network science to the multilayer context, such as centrality measures, among others. This theoretical consolidation sparked extensive research on dynamics across layers, including diffusion, synchronisation, epidemic spreading, and evolutionary games, revealing that interlayer coupling can either enhance or inhibit dynamical processes depending on structural correlations.
In the following years, multilayer methodologies spread rapidly across disciplines. As we will discuss in the following sections, applications now span social, biological, and technological systems. Parallel advances in community detection and inference have made it possible to extract meaningful mesoscale organisation from high-dimensional, heterogeneous data. More recent developments integrate temporal, categorical, and hierarchical aspects within multilayer frameworks and connect them to higher-order network representations such as hypergraphs. Today, the multilayer paradigm stands as one of the most powerful generalizations in network science, bridging theoretical, computational, and empirical approaches to the study of interconnected complex systems.
As such, multilayer network science has become a mature field. In this paper, we provide a high-level overview of progress in this area. We divide the content into two broad parts. In the first, about theory and methods, after a brief introduction of the main concepts, we will review the main techniques to detect communities in multilayer networks, show how dynamic processes unfold on these peculiar networks, summarize progress in the analysis, modelling, and dynamics of temporal networks, explain how higher-order structures can be used to study multilayer networks and how the latter can be projected in vector spaces via neural embeddings. The second part presents the applications of multilayer network science to diverse domains and topics. These include: interdependent systems, spreading processes, computational social science, economic and financial systems, infrastructure and ecological networks, science of science, climate science, network medicine, and network neuroscience. At the end, we highlight current challenges and promising future research directions.
II Theory and Methods
II.1 Fundamental Concepts
II.1.1 Definitions and Formalisms
A network, or graph, is a mathematical structure used to model pairwise relations between objects. It consists of a set of nodes (or vertices) and a set of links (or edges) that connect pairs of nodes. These connections are encoded in the adjacency matrix A, where the element $a_{ij}=1$ if there is a link between nodes $i$ and $j$, and $a_{ij}=0$ otherwise.
A multilayer network generalizes this concept to represent systems that integrate multiple subsystems or types of connectivity. Formally, it is a set of networks, with each network representing a layer of the system. The layers can represent different times, contexts, or types of interaction. Importantly, there can be links within those layers, representing intralayer connections, but also links connecting nodes across layers, interlayer connections.
A prominent example of a multilayer system is the coupling of a power grid and the internet. One layer represents the power grid, with nodes as power stations and substations, while links represent transmission lines, illustrating the distribution of electrical power. Meanwhile, the other layer represents the internet, with nodes as data centres, routing hubs, and servers, linked by data transmission pathways. This setup highlights the interdependency between layers: the internet needs power to function, while modern power grids rely on internet-based communication for their operation, something that can be encoded in the interlayer links. This multilayer network model is essential for studying cascading failures and the overall resilience of interconnected systems Buldyrev et al. (2010); De Domenico (2023).
A multiplex network is a specific and widely used type of multilayer network where nodes represent the same entity on every layer, and interlayer links only connect a node to itself across different layers. Each layer then represents a different type of relationship among the nodes. For instance, a social network can be modelled as a multiplex network where layers represent different platforms or social contexts Kivelä et al. (2014); Battiston et al. (2014).
To analyse multiplex networks, we need to generalize the adjacency matrix. Two primary formalisms are used. First, we have the tensor formalism, in which the separation of layers can be maintained explicitly. A multiplex network can then be represented by an adjacency tensor $\mathcal{M}$ of order 4 with elements $M_{ij}^{\alpha\beta}$. Here, $M_{ij}^{\alpha\alpha}$ represents the weight of the intralayer link between nodes $i$ and $j$ in layer $\alpha$, while $M_{ij}^{\alpha\beta}$ with $\alpha\neq\beta$ represents the interlayer connections. This is a more general and flexible representation, though it requires methods from tensor algebra for analysis De Domenico et al. (2013).
Second, we have the formalism based on the supra-adjacency matrix, which represents the entire multilayer system as a single, larger network. For a multiplex network with $N$ nodes and $L$ layers, the supra-adjacency matrix $\mathcal{A}$ is a $NL\times NL$ block matrix. The diagonal blocks are the adjacency matrices of each layer $\alpha$, $\text{{A}}^{\alpha}$, representing intralayer connections. The off-diagonal blocks encode the interlayer connections, which can be represented by the identity matrix I in the $\alpha\beta$ block if layer $\alpha$ and $\beta$ are connected, and 0 otherwise Kivelä et al. (2014). An example of supra-adjacency matrix of a multiplex network is
$\mathcal{A}=\left(\begin{array}[]{cccc}\text{{A}}^{(1)}&\text{{I}}&0&\cdots\\ \text{{I}}&\text{{A}}^{(2)}&\text{{I}}&\cdots\\ 0&\text{{I}}&\text{{A}}^{(3)}&\cdots\\ \vdots&\vdots&\vdots&\ddots\end{array}\right)$ (1)
This corresponds to a particular case in which only adjacent layers are connected. The advantage of this second formulation is that it is intuitive and allows the direct application of many single-layer network analysis tools to the larger supra-network.
Following these representations, we can also define the corresponding Laplacian matrices. For a single-layer graph, the Laplacian matrix is defined as $\text{{L}}=\text{{D}}-\text{{A}}$, where D is the diagonal matrix of node degrees. Similarly, the supra-Laplacian matrix $\mathcal{L}$ can be constructed from the supra-adjacency matrix $\mathcal{A}$ Kivelä et al. (2014); Gómez et al. (2013). While the supra-adjacency matrix tells us which connections exist and how they are organized, the supra-Laplacian goes a step further; it captures how processes take place along those connections. It extends the analysis from merely identifying the links to understanding how different elements within the network influence each other. This transition from the static structure of connections to the dynamic interplay of influences is pivotal for understanding the overall behaviour of multiplex networks, setting the stage for a deeper exploration of the network’s properties through its eigenvalues Solé-Ribalta et al. (2013).
II.1.2 Structural Metrics
Once a multilayer network is defined, we need a toolkit of metrics to quantify its structure. While many metrics are generalizations from the ones in single-layer networks Aleta et al. (2020b), the most insightful ones are those that capture the unique features arising from the multi-layered structure (see Section II.1.4).
Starting with node-level metrics, the most basic measure of a node’s importance is its degree, $k$, defined as the number of connections it has. In multilayer networks, a node $i$ may have a different number of connections in each layer $\alpha$, yielding the per-layer degree $k_{i}^{\alpha}$. The multi-degree of a node is then the sum of its degrees across all layers. The degree is also a simple measure of the centrality of a node, in that it expresses its importance in the network. Over the last two decades, the list of centrality measures has considerably expanded, and each one of these metrics is defined according to some rules that limit their range of validity and interpretability. Multilayer networks are well-equipped with an extensive list of centrality measures, many of them being generalizations of the corresponding definitions in single-layer networks. Some of them are purely topological, such as the eigenvector, the Katz, and the betweenness centralities De Domenico et al. (2013, 2015b); Solá et al. (2013); Taylor et al. (2021). Others are defined by means of dynamical processes. Notable examples include PageRank Halu et al. (2013) or a centrality related to the random walk occupation probability Solé-Ribalta et al. (2016), among others.
II.1.3 Spectral properties
To characterise the network as a whole, spectral properties of the system’s matrix representations have been widely used. The eigenvalues of the Laplacian matrix of a network offer a wealth of information, which can be utilized to understand various dynamic properties of the network Arenas et al. (2006); Almendral and Díaz-Guilera (2007). The algebraic connectivity or Fiedler value Fiedler (1973), given by the second-smallest eigenvalue ($\lambda_{2}$) of the supra-Laplacian matrix, is a key metric. A larger $\lambda_{2}$ indicates a more cohesive and robust network that is harder to break apart. Conversely, a small $\lambda_{2}$ can indicate potential points of vulnerability or fragmentation within the network. Higher eigenvalues $\{\lambda_{3},\lambda_{4},\,\ldots\}$ offer insights into more complex network structures and dynamicsSolé-Ribalta et al. (2013). For instance, they can be used to detect community structure within the network Chauhan et al. (2009) (see Section II.2.1).
Beyond structural insight, $\lambda_{2}$ plays a fundamental role in dynamical processes such as diffusion, consensus, and especially synchronisation. In networks where each node represents a Kuramoto phase oscillator, $\lambda_{2}$ determines the timescale for convergence to the synchronised state Almendral and Díaz-Guilera (2007). Small values of $\lambda_{2}$ are associated with bottlenecks or weakly connected components that can delay full synchronisation. This spectral framework naturally extends to multiplex networks, where each node is present in multiple layers. Figure 1 shows the theoretical predictions for $\lambda_{2}$ of a two-layer network in the asymptotic regimes: $\lambda_{2}=2D_{x}$ for $D_{x}\ll 1$, and the superposition of both layers for $D_{x}\gg 1$, $D_{x}$ being the interlayer coupling. The plot also includes the eigenvalues of the individual layers and the corresponding supra-Laplacian eigenvalue.
Figure 1: Comparison between the second smallest eigenvalues $\lambda_{2}$ of the different Laplacian matrices as a function of the interlayer coupling $D_{x}$. Reprinted figure with permission from Gómez et al. (2013). ©2013 by the American Physical Society.
Actually, through the behaviour of $\lambda_{2}$ as a function of the interlayer connectivity, we can directly identify when multiplex networks will behave effectively as if each layer were an independent network, or instead when multiplexity kicks in with its emerging effects on the dynamics. This has been shown in a simplified model of a two-layered network where the interlayer connections are weighted by a homogeneous parameter $p$ in Radicchi and Arenas (2013), but the arguments are general for networks with a larger number of layers. The algebraic connectivity of the supra-Laplacian $\mathcal{L}$ reads $\lambda_{2}(\mathcal{L})=2p$ for $p<p^{*}$ and is upper-bounded such that $\lambda_{2}(\mathcal{L})\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/2$ for $p\geq p^{*}$, the critical point being $p^{*}\leq\lambda_{2}(\mathcal{L_{A}}+\mathcal{L_{B}})/4$. As it can be guessed graphically, but proved mathematically, the first derivative of $\lambda_{2}(\mathcal{L})$ is discontinuous. This emerges as a result of the crossing of two distinct groups of eigenvalues. Actually, these crossing events appear for the other eigenpairs of the graph Laplacian, with the exception of the smallest and largest eigenvalues. Consequently, this results in discontinuities in the first derivatives of the corresponding eigenvalues Radicchi and Arenas (2013).
The stability of the synchronised state in these systems can be analysed using the Master Stability Function (MSF) formalism Pecora and Carroll (1998), which allows us to reduce the problem to a spectral one. The synchronisability of a network can be assessed by the eigenratio
$R=\frac{\lambda_{N}}{\lambda_{2}},$ (2)
where $\lambda_{N}$ is the largest eigenvalue of the Laplacian matrix. A smaller eigenratio $R$ implies greater synchronisability, since it indicates a wider range of coupling strengths for which the synchronised solution remains stable.
In Solé-Ribalta et al. (2013), analytical approximations for the eigenratio $R$ were provided in two limiting regimes: weak interlayer coupling ($D_{x}\ll 1$) and strong interlayer coupling ($D_{x}\gg 1$). These approximations offer explicit insight into how the topology and interlayer coupling influence the dynamics. The predictions were shown to agree well with numerical results in a wide range of coupling values, as illustrated in Figure 2. The point at which the two analytical regimes intersect provides an estimate of the optimal coupling strength for synchronisation performance.
Figure 2: Eigenratio $R=\lambda_{N}/\lambda_{2}$ as a function of the interlayer coupling strength $D_{x}$. Circles represent exact numerical results, while the solid lines show the analytical approximations valid in the weak- and strong-coupling limits. Adapted figure with permission from Solé-Ribalta et al. (2013). ©2013 by the American Physical Society.
These results support the use of spectral indicators, particularly $\lambda_{2}$ and the eigenratio $R$, as powerful tools for connecting structure and dynamics in multilayer networks.
II.1.4 Multilayer metrics
Some metrics and concepts do not have counterparts in standard single-layer networks. The characterisation of multilayer networks through these new metrics has unveiled an array of phenomena hitherto unobserved on single-layered systems, thus manifesting the singularity and usefulness of the multilayer description. For instance, the participation coefficient measures how evenly a node’s links are spread across the different layers. A node with high multi-degree but low participation coefficient is a hub - a highly connected node - in only one or a few layers, while a node with high participation coefficient is genuinely active across the entire system. This distinction is crucial for understanding a node’s role in processes that span multiple layers Battiston et al. (2014).
Another network-level metric with no single-layer equivalent is the mutually largest connected component (MLCC). This is the largest group of nodes where every node can reach every other node through paths existing within each and every layer Buldyrev et al. (2010). The size of the MLCC is a powerful indicator of a system’s integrated resilience Artime et al. (2024). In many multilayer systems, reducing network connectivity (e.g., by removing nodes or links) can cause the MLCC to collapse abruptly in a discontinuous phase transition, contrasting sharply with the smoother, continuous transitions typically seen in single-layer networks Gao et al. (2012); Radicchi and Arenas (2013)
A fundamental question in multilayer network science is: “What new information does the layered structure provide compared to a simple, aggregated network?” To answer this, we need metrics that quantify the relationships between layers. One approach is to measure layer similarity or redundancy. This can be done by calculating the interlayer mutual information or other correlation measures between the adjacency matrices of different layers. If layers are highly similar, the system is redundant, and an aggregated view might suffice. If they are diverse, the multilayer representation is essential, as different layers capture distinct relational patterns De Domenico et al. (2015b); Schieber et al. (2017); Ghavasieh and De Domenico (2020).
Furthermore, multilayer networks introduce new types of structural correlations. Beyond intralayer patterns like degree assortativity (whether high-degree nodes connect to other high-degree nodes Pastor-Satorras et al. (2001); Newman (2002)), we can measure interlayer correlations. For example, are nodes with a high degree in one layer also likely to have a high degree in another? The presence or absence of such correlations significantly impacts the behaviour of dynamical processes, from spreading to synchronisation, and has been a major focus of research Pamfil et al. (2019); Nicosia and Latora (2015); Wu et al. (2020); Nicosia et al. (2013a); Bianconi (2013); Battiston et al. (2014); Menichetti et al. (2014); Lee et al. (2012).
II.2 Structure and Dynamics
II.2.1 Community detection
How shall we define communities in multilayer networks? The answer is context-dependent. In a network of customers layered by product categories, a community might be a group of people with consistently similar purchasing habits across multiple layers (a cross-market segment). In a biomedical context, coupling a gene co-expression network with a protein-protein interaction network, a community could reveal groups of genes that are co-expressed and whose encoded proteins physically interact. The search for such meaningful groups in multilayer networks has become a vibrant research area, with several reviews already charting the landscape Kim and Lee (2015); Magnani et al. (2021a); Huang et al. (2021). Here, we summarize the main strategies and highlight key challenges and future directions.
Community detection is mostly performed on multiplex networks, where the same node exists across all layers. This framework is also directly applicable to temporal (multi-slice) networks, treating each time snapshot as a layer (see Section II.3.1). Methodologically, approaches to find multilayer communities can be grouped into three main strategies: flattening, layer-by-layer, and multilayer.
The first strategy, flattening, collapses the multilayer network into a single aggregated graph containing all nodes and links from all layers Berlingerio et al. (2011); Kim et al. (2016). Any traditional algorithm working on single-layer networks can then be applied. While computationally efficient, this approach discards all information about the distinct layers and their relationships.
Layer-by-layer algorithms first detect communities on each layer independently and then aggregate the results. These are further divided into three classes: frequent subgraph mining Zeng et al. (2006); Berlingerio et al. (2013), feature integration Dong et al. (2013); Nie et al. (2018), and partition integration Cantini et al. (2015a); Tagarelli et al. (2017). Frequent subgraph mining is based on the idea that the nodes of a multilayer community are assigned to the same group in different layers. Thus, it relies on identifying groups of nodes that are consistently clustered together across many layers. Feature integration relies on the principle that communities are made of nodes that are similar to each other, something typically measured by looking at the eigenvectors of the layers. Communities are found by grouping nodes according to their features via, e.g., k-means clustering Forgy (1965); MacQueen (1967). Partition integration directly combines the divisions found on the individual layers. For instance, one could derive a consensus partition, which is the division that is closest to all the input ones, on average Strehl and Ghosh (2002).
Figure 3: Multilayer communities. The continuous (red) contours indicate communities. (a) Intralayer communities group state nodes on the same layer. (b) Members of cross-layer communities can be on both the same and different layers. (c) Pillar communities group all the counterparts of the nodes in other layers in the same community.
Lastly, we have genuinely multilayer methods. These algorithms operate on the complete multilayer structure without prior aggregation. Furthermore, they do not cluster single nodes but state nodes (or node-layer tuples, i.e., a specific node on a specific layer). This allows for richer community definitions (Fig. 3):
  • Intralayer communities group state nodes on the same layer.
  • Cross-layer communities can group state nodes from different layers.
  • Pillar communities are a constrained type where if a node belongs to a community, all of its counterparts in other layers also belong to the same community.
Many of these methods are a direct extension of single-layer algorithms to the multilayer case. Modularity maximization was generalized by defining a multilayer null model based on random walks, with intralinks and interlinks playing different roles Mucha et al. (2010). Local exploration methods, like seed set expansion Interdonato et al. (2017), label propagation Boutemine and Bouguessa (2017) or the clique percolation method Afsarmanesh Tehrani and Magnani (2018), as well as random walk-based algorithms such as Infomap De Domenico et al. (2015a) and Walktrap Kuncheva and Montana (2015), have also been adapted. Furthermore, statistical inference methods, primarily based on the stochastic block model (SBM), offer a principled framework to detect communities, infer relationships between layers, and even test whether an observed single-layer network is better explained as an aggregation of hidden layers Peixoto (2015); De Bacco et al. (2017).
Despite this progress, significant challenges remain. First, a persistent issue is the rigorous assessment of statistical significance. Without it, one cannot distinguish meaningful communities from patterns arising by chance. Second, the increasing availability of node and edge metadata is often used only for validating results, a procedure known to be unreliable Hric et al. (2014); Peel et al. (2017). Instead, this information could be directly integrated into the detection process to find more accurate and relevant communities, a task for which SBM-based approaches are particularly well-suited Hric et al. (2016); Newman and Clauset (2016). Finally, the vast majority of methods are designed for multiplex networks. There is a clear need for algorithms that can handle more general multilayer networks (often called heterogeneous networks in computer science) with different node sets and relationships in each layer, a promising direction for future research Sun and Han (2013).
II.2.2 Dynamics and Processes
The multiplex network framework allows for a wide variety of dynamical behaviours that go beyond simple extensions of single-layer dynamics. Unlike isolated networks, where each node interacts through a single set of connections, multiplex networks allow each node to participate in multiple interaction layers, each possibly governed by different dynamics, rules, or time scales. This structure enables novel emergent behaviours that are fundamentally rooted in the interplay between dynamics on and across layers.
In some fields, particularly computer science, it is common to distinguish between actors (the underlying real-world entities) and nodes (their representations within specific layers) Czuba et al. (2024); Zhong et al. (2022); Czuba et al. (2025). While this specific terminology may be less explicit in other domains, the underlying principle of defining processes at the level of the entity, rather than solely at the layer-specific node, is widely relevant. This framework allows dynamics to be driven by the actor, integrating states from all its corresponding nodes simultaneously, or by specific layer-level copies. Consequently, it enables complex activation rules: an actor might change state only if a threshold of its nodes are active (reinforcement), or conversely, activity in a single node might drive the entire actor (contagion).
In what follows, we review several representative classes of dynamical processes that have been studied in the context of multiplex networks, emphasizing those in which the multilayer structure is essential to the observed phenomena. The discipline-specific applications of these processes will be presented later in Section III.
A paradigmatic case in which multiplexity alters the behaviour of a well-understood process is that of reaction-diffusion systems. In multiplex networks, reaction and diffusion terms can be assigned to different layers or coupled across them. For instance, in Kouvaris et al. (2015), activator and inhibitor species are placed on separate layers, where each species diffuses only within its own layer but reacts across layers. Remarkably, the presence of interlayer coupling allows for Turing-like pattern formation even when both species have identical mobility rates—something impossible in single-layer networks. This form of topology-driven instability, further explored in Asllani et al. (2014), highlights how multiplexity can give rise to heterogeneous and robust spatial patterns under conditions where the dynamics of isolated layers remain homogeneous.
In multiplex systems, one well-studied example of interacting dynamics is the coupling between epidemic spreading and awareness diffusion Granell et al. (2013). In this model, the disease propagates through physical contact on one layer, while information about the disease (awareness) spreads via virtual communication on a separate layer. The two processes mutually inhibit each other, giving rise to a point at which the presence of awareness fundamentally alters both the epidemic threshold and the eventual prevalence.
Multiplex networks also provide a natural setting for studying feedback between processes on different layers. In Nicosia et al. (2017), a model is introduced in which neural synchronisation on one layer modulates resource allocation on another, and vice versa. This mutual feedback leads to the spontaneous emergence of explosive synchronisation and highly heterogeneous resource distributions.
Multiplex networks often feature layers with different topologies or dynamics, leading to hybrid systems. In Hizanidis et al. (2016), a modular neural network is modelled with electrical synapses within communities and chemical synapses across them, each type represented as a different layer. The interplay between these types of couplings gives rise to chimera-like states, where coherent and incoherent activity coexist across modules. Such hybrid multiplex models are relevant to understanding how local synchrony coexists with global disorder in brain dynamics.
A similar approach is taken in Sadilek and Thurner (2015), where a two-layer Kuramoto model is derived from physiologically motivated equations for cortical dynamics. The model captures transitions between resting and background brain activity, successfully reproducing empirical features that were inaccessible to simpler models. Again, multiplexity allows for distinct coupling schemes across layers (e.g., electrical vs chemical, or fast vs slow processes) that shape the global phase diagram.
Another work examined how two layers not directly connected can synchronise through the mediation of a central relay layer Leyva et al. (2018). Results showed that synchronisation thresholds are reduced in the multiplex setting and that low-degree nodes play a crucial role, while hubs can be “demultiplexed” without disrupting global coherence. More recently, a three-layer multiplex of non-linear oscillators, in which a central drive layer unidirectionally couples to two response layers, has been investigated Vadakkan and Ambika (2025). Such directional coupling induces synchronised oscillations in the response layers, even with amplification relative to the drive, and varying interlayer strength or introducing time-scale mismatches enables control over synchronisation patterns, amplitudes, and dynamical regimes such as quasi-periodicity and generalized synchronisation.
These examples collectively illustrate that multiplex networks enable a new class of dynamical processes: systems where different types of dynamics not only coexist but interact, often through cross-layer mechanisms. This interplay can give rise to topology-induced instabilities, extended metastable states, abrupt phase transitions, novel critical points, and even the coexistence of order and disorder.
Such behaviours are not mere quantitative variations of single-layer dynamics but emerge qualitatively due to the multiplex architecture. Importantly, these phenomena are often robust to parameter variations and are observed across different domains, from neuroscience and epidemiology to social systems. In Section III, we will revisit some of these dynamical mechanisms in the context of concrete applications, where domain-specific constraints and empirical data can further shape the interaction between structure and dynamics.
II.3 Advanced Frameworks
II.3.1 Temporal Networks
Complex systems, from social groups to interacting neurons, are fundamentally dynamic. While classic network science often represents these systems with static graphs, this is an aggregation that loses all temporal information, such as the sequence, duration, or concurrency of interactions. The recent availability of time-resolved data and computational power has propelled the study of temporal networks, revealing phenomena invisible in a static view Holme and Saramäki (2012); Holme (2015). A temporal network can be represented as a sequence of static snapshots, which naturally fits the multilayer network framework where each layer is a snapshot in time. Besides, it is also possible to have networks that are temporal and multilayer at the same time, in which case the temporal correlations between layers can significantly affect the dynamics of processes unfolding on multilayer networks Starnini et al. (2017).
While early theoretical models added temporality by dynamically changing links, particularly in an adaptive fashion Gross and Blasius (2008), the strong expansion of the field has been mostly data-driven Holme and Saramäki (2012); Holme (2015). The availability of high-resolution temporal data, such as the face-to-face contact network in a museum exhibition depicted in Figure 4 and collected by the SocioPatterns collaboration using wearable sensors Cattuto et al. (2010), provided a strong empirical foundation for the field. This data made it clear that static representations not only hide temporal details but can be misleading, as the properties of a networked system can vary dramatically from one moment to the next Masuda and Lambiotte (2016); Saramäki and Moro (2015).
Figure 4: Temporal networks of face-to-face interactions of people on different days of museum visits. In each day, time evolution is represented as a colour code. The data has been recorded by detecting physical proximity using wearable sensors developed by the SocioPatterns collaboration Isella et al. (2011); Van den Broeck et al. (2012). Reprinted figure with permission from the SocioPatterns collaboration.
This data-centric view is what led researchers to analyse temporal networks as a series of snapshots. This approach revealed that while aggregated properties might converge to the static ones Krings et al. (2012), local measures like a node’s centrality could fluctuate significantly, with nodes shifting from central to peripheral roles over time Braha and Bar-Yam (2006, 2009). However, a true paradigm shift occurred when the focus moved from fluctuating static properties to inherently temporal concepts. Researchers recognized that information flows along temporal paths, which are a set of time-respecting sequences of interactions Kempe et al. (2000); Holme (2005). A path that exists in an aggregated static network may be impossible in a temporal one if the interactions do not occur in the correct chronological order Pan and Saramäki (2011). This insight spurred the redefinition of classical centrality measures for the temporal domain Stehlé et al. (2010) and led to the discovery of new phenomena, such as the burstiness of human interactions, where events are clustered in time rather than occurring randomly Barabasi (2005); Karsai et al. (2018).
But not all temporal interactions are equally informative. Patterns may appear briefly, recur, or vanish entirely. Motifs that are evident in aggregated static views of the network may never exist in real time due to improper chronological alignment Holme (2005). To deal with this complexity, researchers have developed various techniques for filtering and pattern detection. At the local scale, temporal backbones highlight node pairs whose interaction frequencies significantly exceed what would be expected under a temporal null model Kobayashi et al. (2019). Temporal motifs identify frequently recurring, time-ordered subgraphs Kovanen et al. (2013), while temporal cores generalize the static core decomposition to uncover dense substructures active over defined time windows Galimberti et al. (2018). Similarly, the temporal rich-club coefficient measures whether highly connected nodes in the aggregate network tend to form simultaneous, persistent connections Pedreschi et al. (2022).
In parallel, significant effort has gone to identifying the relevant timescales that govern the evolution of temporal networks. Early techniques relied on comparing successive time windows Darst et al. (2016), but this overlooked long-range correlations. More sophisticated methods compute pairwise distances or similarities between network snapshots Masuda and Holme (2019), enabling the clustering of recurring states and the identification of their transitions. Other approaches apply auto-correlation metrics, or Fourier transforms to extract temporal correlations and recurrent interaction patterns at different time scales Lacasa et al. (2022); Andres et al. (2024a).
Representing temporal networks in a simple form for analysis has also been a priority. Two lossless, static-like representations have emerged as especially useful. The supra-adjacency representation encodes node-time pairs as nodes in a multilayer structure, linking them according to the observed interactions Valdano et al. (2015); Sato et al. (2021). The event graph, in contrast, treats interactions themselves as nodes, linking temporally adjacent events involving shared participants Kivelä et al. (2018). This latter framework is especially well-suited to model processes constrained by timing, such as epidemic spreading or transit dynamics Badie-Modiri et al. (2022). Another line of work uses graph embedding techniques to map high-dimensional temporal networks into low-dimensional vector spaces Goyal and Ferrara (2018); Cai et al. (2018). These embeddings preserve structural or temporal similarities and support tasks such as link prediction or influence sets Thongprayoon et al. (2023); Torricelli et al. (2020); Dall’Amico et al. (2024). Embeddings also facilitate data compression and learning-based applications, as discussed further in Section II.3.3.
Models of temporal networks have been developed to generate realistic synthetic networks or to establish null benchmarks. Mechanistic models aim to reproduce observed features or test hypotheses about dynamics Holme and Liljeros (2015), while randomized models offer a controlled way to destroy correlations and assess their significance Gauvin et al. (2022). These models range from memoryless activity-driven frameworks Perra et al. (2012); Starnini and Pastor-Satorras (2013) to non-Markovian extensions capturing burstiness and memory effects Moinet et al. (2015). Additional approaches include stochastic block models, priority queue models, models based on Hawkes or other dynamical processes, and memory network models.
The temporal networks framework also adds another layer of complexity to dynamical processes unfolding on them, since one needs to take into account the coupling between the characteristic timescale of the dynamic process and the temporal network’s one. In epidemiology, the coupling between interactions and disease progression profoundly affects spreading outcomes with important consequences for designing effective public health interventions Masuda and Holme (2017); Machens et al. (2013); Koher et al. (2019); Toroczkai and Guclu (2007); Valdano et al. (2015). In social systems, temporal networks are used to model the spread of information, behaviours, and innovations, often through threshold mechanisms distinct from disease transmission Unicomb et al. (2021); Andres et al. (2024b). Temporal opinion dynamics have been explored to understand consensus and polarization Li and Dankowicz (2019); Chu and Porter (2023), while synchronisation phenomena in biology and engineering are modelled as coupled oscillators interacting over time-varying links Ghosh et al. (2022). Applications in transportation include traffic optimisation and infrastructure design Salama et al. (2022).
Looking forward, a central challenge is identifying the minimal yet sufficient temporal structures necessary for specific applications. For example, while the temporal paths in a physical proximity network may code the ways an epidemic would spread between the people, several interactions may not fall on these paths and thus appear completely irrelevant from the aspect of a spreading process. Thus, since many interactions may not influence a process, reducing networks to effective backbones could lead to smaller and faster computations, as well as improving data sharing in a privacy-preserving way.
II.3.2 Higher-order interactions
Group interactions are widespread in real-world systems, ranging from social dynamics Iacopini et al. (2019); de Arruda, Guilherme Ferraz et al. (2020); Battiston et al. (2021); Landry and Restrepo (2020); St-Onge et al. (2022); Ferraz de Arruda et al. (2023); Kiss et al. (2023), such as group chats, to chemical reactions involving multiple reactants Jost and Mulas (2019); Battiston et al. (2021); Traversa et al. (2023). In these contexts, the traditional network science framework, which considers only pairwise interactions, is often insufficient to capture the complex dependencies between multiple interacting entities. Higher-order models provide a richer framework for accurately representing these systems and avoiding misleading predictions Battiston et al. (2020, 2021); Ferraz de Arruda et al. (2023, 2024). For example, contagion processes that model the population as a graph will present a second-order phase transition Pastor-Satorras et al. (2015); de Arruda et al. (2018). However, contagion models in higher-order networks often exhibit discontinuous transitions Iacopini et al. (2019); de Arruda, Guilherme Ferraz et al. (2020); Landry and Restrepo (2020); St-Onge et al. (2022); Ferraz de Arruda et al. (2023, 2024). Similarly, in pairwise contagion models, one expects a single stable solution, while multistability is possible in contagion models in higher-order networks Ferraz de Arruda et al. (2023); Kiss et al. (2023); Ferraz de Arruda et al. (2024). Interestingly, these new phenomena are not restricted to contagion processes; they have also been observed in synchronisation processes Skardal et al. (2023) and evolutionary games Civilini et al. (2024). These examples demonstrate how neglecting the higher-order nature of a process can alter the outcome of a model, potentially leading to incorrect predictions.
Mathematically, higher-order interactions are usually modelled using hypergraphs or simplicial complexes Battiston et al. (2020); Boccaletti et al. (2023); Torres et al. (2021). Hypergraphs allow edges (hyperedges) to connect any number of nodes, whereas simplicial complexes require that all subsets of a hyperedge be present. The best choice of representation often depends on the dynamics being studied. A discussion about this topic can be found in Torres et al. (2021).
Figure 5: Schematic representation of different networked systems. The bottom row depicts systems with pairwise interactions, while the top row contains group interactions including more than two nodes. As with classical pairwise models, it is possible to associate a type to each interaction, extending the framework of multilayer networks to multilayer hypergraphs. Adapted figure with permission from Ferraz de Arruda et al. (2024). ©2024 by Springer Nature.
Compared to pairwise networks, the structural and spectral analysis of higher-order systems is still in development. While concepts such as adjacency and Laplacians are well-defined in networks, their hypergraph counterparts are more complex and often combine structural and dynamical parameters Schaub et al. (2020); Mulas et al. (2020); Ferraz de Arruda et al. (2021). This complicates the analytical study of processes such as diffusion and synchronisation, and different definitions may yield different outcomes depending on the context. Similarly, most inference methods in the literature deal with inferring pairwise relationships. Detecting higher-order structures requires new methods Young et al. (2021); Contisciani et al. (2022).
Recent work has begun to explore multilayer extensions of higher-order systems (Fig. 5), such as triadic percolation processes and interdependent dynamics across layers Sun et al. (2023); Chang et al. (2023). In the context of spreading processes, a two-layer approach considering the concurrent spread of information and disease was studied in Fan et al. (2022); Hong et al. (2023). Furthermore, the diffusion of resources coupled with epidemic spreading was examined in  Sun et al. (2022). These studies demonstrate how combining multilayer and higher-order structures can lead to novel, emergent behaviours, such as oscillatory or chaotic dynamics. Another promising research direction is studying temporal hypergraphs, which extend these representations to time-varying settings. This framework captures interactions involving groups of nodes that occur over finite durations. Examples include simultaneous face-to-face contacts Di Gaetano et al. (2024); Gallo et al. (2024a) and group-level dynamics in social settings Iacopini et al. (2024). Temporal hypergraphs provide a natural language to describe and model these evolving group structures, opening new avenues for analysing their evolution and how they influence dynamical processes.
II.3.3 Machine learning
Graph representation learning, or graph embedding, aims to map nodes into a geometric space such that structural properties are preserved Goyal and Ferrara (2018); Cai et al. (2018). This task is essential for applying machine learning techniques to relational data and has been widely used for link prediction, node classification, routing, and visualisation. In the past few years, traditional embedding methods have been extended to deal with heterogeneous graphs and knowledge graphs, where nodes or edges carry additional categorical, numerical or textual information Shi et al. (2022). Multiplex networks can be seen as a specific type of heterogeneous graph and have prompted the development of dedicated embedding techniques.
A good multiplex embedding method should effectively capture both intralayer and interlayer structures. Multiplex network embeddings can be divided into two major classes: joint embeddings, which project all replicas of a node across layers to a single point in a low-dimensional Euclidean space, and individual embeddings, where each layer has its own representation of the node. Joint embeddings can be obtained by merging individual embeddings. The general approach followed by most multiplex embedding techniques consists of three phases: sampling, training, and optimisation.
Figure 6: General approach to multiplex embedding. Given a multiplex network, the first step involves identifying node pairs that are supposed to have similar (positive samples) and dissimilar (negative samples) representations. Next, the learning framework is used to train the node representations based on the sampled data. The node representations can be learnt for the entire multiplex (bottom right) or for individual layers (top right). Individual embeddings may be fused together to obtain joint embeddings.
Sampling. While embedding, we want pairs of nodes that are close to each other in the network to also have similar embeddings. These are called “positive node pairs” and can be found using several strategies. Techniques such as DeepWalk Perozzi et al. (2014) and node2vec Grover and Leskovec (2016) simulate random walks to define contexts, akin to word embeddings in natural language processing Mikolov (2013). These approaches have been extended to multiplex networks by allowing walks to transition between layers with some probability Liu et al. (2017). Other methods define proximity using K-nearest neighbors or neighborhood-based criteria, as in GraphSAGE Hamilton et al. (2017). To reduce computational cost, negative sampling is used to approximate objective functions, drawing random “negative pairs” assumed to be dissimilar Cai et al. (2018).
Training. In this phase, multiplex embedding methods learn representations using shallow neural networks, although more advanced architectures have been proposed. In shallow neural networks, the training process is specified by the structural organisation of the learning algorithm, which can be layerwise or collective learning (Fig. 6, top and bottom rows respectively). Layerwise learning separately trains node embeddings for each layer and combines them using weights, often learned via attention mechanisms Qu et al. (2017); Xie et al. (2020). Collective learning directly learns unified representations by leveraging multiplex sampling and skipgram-style optimisations Zhang et al. (2018); Liu et al. (2017). While shallow networks are common, they face important limitations, such as the inability to generalize to unseen nodes or share parameters between nodes (since the encoder optimises a unique embedding vector for each node). To alleviate these limitations, more sophisticated learning architectures, like graph neural networks, have been proposed.
Optimisation. Embeddings are adjusted to minimize loss functions that reward proximity between positive pairs and penalize proximity between negative ones. Some methods also include node attributes to improve representation quality Feng et al. (2019); Cen et al. (2019). These optimised embeddings can then be used in downstream tasks such as:
  • Link prediction: predicting missing or future links within a layer (layer-wise), across layers (interlayer) or in an aggregated view of the multilayer Liu et al. (2017); Xie et al. (2020); Zhang et al. (2018); Feng et al. (2019); Gallo et al. (2024b).
  • Node classification: predicting labels for unlabelled nodes based on known node attributes or partial label data Zhang et al. (2018).
  • Community detection: discovering modular structures across layers by learning weights that reflect each layer’s contribution Jing et al. (2021); Cai et al. (2021).
In addition to Euclidean embeddings, hyperbolic embeddings provide a geometry better suited to hierarchical and navigable networks García-Pérez et al. (2019). These approaches can potentially improve performance in tasks like greedy routing, where shortest paths are found using only local information. Recent work using layerwise hyperbolic embeddings suggests that interlayer geometric correlations improve routing performance Kleineberg et al. (2016). Future work could explore hyperbolic multiplex embeddings more systematically.
Lastly, synthetic multiplex models Bazzi et al. (2020) can be used to systematically vary layer-specific properties and interlayer correlations, providing testbeds for evaluating embedding performance under controlled conditions. These models support the study of how properties like modularity and community alignment across layers affect embedding quality and task performance.
III Applications
Network science has long served as a unifying language across disciplines, offering a flexible framework for modelling complex systems. With the rise of multilayer network theory, this versatility has only expanded. Multilayer networks allow researchers to capture the rich, interconnected structure of systems where multiple types of interactions, temporal dynamics, or functional interdependencies coexist. As a result, their applications span a diverse array of domains—from infrastructure and biology to economics and the social sciences—demonstrating both the theoretical depth and real-world impact of the multilayer perspective.
III.1 Interdependent systems
The importance of multilayer interdependent networks was first recognized by engineers and later formalized through the interdependent percolation model introduced in Buldyrev et al. (2010). In addition to standard intralayer links, these systems feature dependency links that connect nodes across layers, such that the functionality of a node in one network depends on the functionality of a node in another. This structure can lead to cascading failures as a breakdown in one layer propagates to others, potentially causing large-scale system collapse. Unlike standard percolation, where failure progresses gradually, interdependent networks can experience abrupt, catastrophic transitions. Notably, networks with heterogeneous degree distributions, typically robust against random failures, become vulnerable in the presence of dependency links, highlighting important design and resilience considerations.
Insights into these phenomena have been obtained through both analytical and numerical studies Gao et al. (2012); Havlin et al. (2015); Radicchi (2015). In interdependent random networks, the percolation transition is of a hybrid nature: it is discontinuous but also exhibits critical phenomena, such as the presence of critical exponents Baxter et al. (2012). In plain percolation, there are two equivalent theoretical descriptions, one based on cluster statistics and the other on the order parameter Stauffer and Aharony (2018). Yet, these approaches split in the interdependent case, leading to two sets of critical exponents related by scaling relations Lee et al. (2016); Choi et al. (2024). In a recent paper, an interesting relationship between the hybrid transition in cascading processes and the spinodal one was shown, leading to the identification of universality classes Bonamassa et al. (2025).
One of the most exciting recent developments is the experimental realization of interdependent percolation in coupled superconducting multilayer systems Bonamassa et al. (2023). In this setup, two layers of disordered superconductors are separated by an insulating layer but become effectively coupled through Joule heat. This way, when one superconducting grain transitions to a normal state, it generates heat that causes other grains in the adjacent layer to transition as well. The resulting feedback loop drives a non-local dependency coupling and confirms the theoretical prediction of an abrupt transition. These experimental observations not only validate the theory but also open the door to innovative technological applications based on interdependent dynamics.
III.2 Spreading processes
In the context of diffusion and spreading dynamics, multilayer networks offer a powerful framework to study how coupled processes may interact, leading to the uncovering of hidden dependencies, feedback loops, and emergent phenomena de Arruda et al. (2018). This approach is particularly valuable when different types of interactions or dynamical processes influence each other in ways that are not evident in single-layer representations.
The range of systems that benefit from this approach is broad. For example, the mutual influence between public opinion and disease spread can be modelled by layering the dynamics of information diffusion with those of epidemic propagation. Such an approach reveals insights into behavioural responses and feedback mechanisms that would otherwise be obscured Granell et al. (2013); Wang et al. (2015); Fosch et al. (2023); Peng et al. (2021). Multilayer networks also naturally represent systems involving multiple domains with interdependent effects, such as the bidirectional impact between pandemics and economic activity (Fig. 7Pangallo et al. (2024). Or systems with a single process but with different spreading properties depending on the type of interaction Aleta et al. (2020a).
Another compelling example involves the interplay between diseases with distinct transmission mechanisms. Tuberculosis (TB), transmitted via airborne particles, and HIV, spread through sexual contact, represent independent but coexisting threats within a population. Using a multilayer framework, one can explore how interventions or behavioural changes in one disease layer affect the dynamics of the other Sanz et al. (2014); Bosetti et al. (2020); Kinsley et al. (2020); Wang et al. (2019). This holistic modelling approach aids in designing integrated public health strategies.
In essence, multilayer networks provide a powerful framework to address the intricate interdependencies between different spreading processes, opening new avenues for understanding, forecasting, and controlling these complex phenomena.
Figure 7: Schematic representation of an epidemic-economic model. The figure illustrates the interplay between two very different dynamics. An economic component, which represents the flow of goods and services among industries and to final consumers through an input-output network, and an epidemic component, tracking pathogen exposure in various settings such as workplaces, community/consumption venues, schools, and households via a multilayer contact network. Agents within the system exhibit diverse socio-economic characteristics (see the accompanying box). The tight integration of the economic and epidemic modules is evident: the epidemic module influences economic outcomes by reducing consumption due to infection concerns, while the economic module affects epidemic transmission by modifying workplace and community interactions through industry-specific employment fluctuations. Reprinted figure with permission from Pangallo et al. (2024). ©2024 by Springer Nature.
III.3 Computational social science
Social systems are inherently characterised by multiple types of relationships that coexist and interact across diverse contexts, timescales, and communication channels. In computational social science, multilayer networks have become a key framework for analysing how behaviours, attitudes, and information propagate through these interconnected pathways. This approach enables the study of complex dynamics such as the interplay between online and offline ties Hristova et al. (2014), cross-platform relationship evolution Ureña-Carrión et al. (2024), and behaviours emerging from interacting relationship types Garcia et al. (2015).
In political communication, distinct interaction types (support, comments, likes) form separate layers of a multiplex network, each reflecting unique dimensions of polarization and temporal evolution Garcia et al. (2015). For example, in parliamentary Twitter (now X) networks, echo chambers differ across layers: following networks are highly polarized, while mention networks allow more cross-ideological communication Del Valle and Bravo (2018). Cross-platform studies find consistent drivers of polarization, such as homophily and selective exposure, even when patterns vary between platforms Cinelli et al. (2021). In a study of French Twitter users, the multidimensional character of polarization was revealed by unfolding the interaction between the layer of professional politicians and that of ordinary users Peralta et al. (2024).
Temporal multilayer networks reveal that social multiplexity itself is dynamic. Relationship types activate in different contexts, and temporal correlations between layers modify how influence and information spread Ureña-Carrión et al. (2024). Human activity’s bursty nature can create both real and spurious correlations Starnini et al. (2017). In signed networks Diaz-Diaz et al. (2025), research has explored how structural balance and polarization evolve, including dynamical equilibrium where macro-level patterns persist despite micro-level changes Talaga et al. (2023); González-Casado et al. (2025).
In the semantic domain, multilayer networks help uncover how communities structure discourse around complex issues. Modelling words, themes, and temporal evolution as distinct layers reveals patterns of resilience and vulnerability in discussions, such as those following disasters Randazzo et al. (2024). Temporal text networks Vega and Magnani (2018) and multimodal network models Bonifazi et al. (2023) further support analysis of communication across media and platforms.
Multilayer approaches have also enriched opinion dynamics, where single-layer models often oversimplify how ideas form and spread. This approach provides a powerful framework for exploring how phenomena in one domain can influence behaviour in another. For example, a model using a communication layer for sharing opinions and an influence layer for observing actions, revealed a range of complex real-world outcomes Zino et al. (2020). These included rapid, widespread paradigm shifts, the establishment of unpopular norms where innovation adoption fails despite positive sentiment, and persistent community resistance to change.
As discussed in the previous section, the interaction between different processes is a key area of study for multilayer network research. A recent work Amato et al. (2017) investigated a two-layered network, where agents held opinions on different issues in each layer. While each layer individually tended toward consensus, the interactions between layers could stabilize a mixed state where both opinions coexisted for a long period. This was most pronounced in sparse, positively correlated networks, which facilitated the formation of cross-layer opinion groups.
The adaptive voter model Holme and Newman (2006) has been extended to multilayer networks to study how evolving connections affect collective opinion. While multiplexity (the fraction of common nodes across layers) can prevent complete fragmentation, it can also give rise to a new “shattered” fragmentation phase Diakonova et al. (2016). This phase is characterised by large consensus groups of opposite opinions coexisting with isolated nodes. A similar shattered fragmentation transition in the multiplex adaptive voter model with a notion of triadic closure has been used to describe the empirical distribution of cluster sizes in an online game Klimek et al. (2016b).
Furthermore, a central question in this modelling paradigm is the extent to which such complex multilayer systems can be simplified, especially whether the dynamics on a multilayer network could be reduced to an equivalent single-layer representation. It turns out that conventional aggregation procedures fail to capture key non-linear effects, such as the prolonged lifetime of multiplex states at low interlayer connectivity Diakonova et al. (2016). This work suggests that there are fundamental, non-reducible differences between multilayer dynamics and their single-layer approximations, highlighting the necessity of explicitly modelling these intricate, interconnected structures to fully understand complex social phenomena.
Altogether, the integration of multilayer networks into computational social science enables more realistic, nuanced models of human behaviour. As interactions increasingly span digital and physical spaces, these tools are essential for understanding polarization, social cohesion, and collective action in a complex, connected world.
III.4 Economic and financial systems
Economics and finance are inherently based on networked systems. They describe interactions among diverse actors, including individuals, firms, governments, and institutions, that trade, collaborate, invest in, or oversee one another. Markets can be seen as platforms that promote the constant evolution and reshaping of existing networks among different actors. For instance, financial markets facilitate the creation and deletion of links between investors and firms, while the labour market connects workers with employers. New technological platforms can also rapidly reshape existing networks, as in the case of Uber transforming urban transportation labour markets. In all these cases, economic and financial systems heavily rely on information flowing between entities belonging to different networks, such as price signals, market trends, political news, and policy decisions.
Based on this premise, it is not surprising that multiplex and multilayer networks may occupy a central position in the analysis and understanding of economic and financial systems. While challenges such as data access and system complexity have limited the full adoption of these approaches, notable studies already highlight their potential. One example is the use of a multiplex global financial network, in which nodes represent countries, edges encode cross-country financial assets of various types, and layers represent asset types. The number of systemically important countries almost doubles when the heterogeneity of financial exposures is considered (i.e., when using a multiplex network), compared to results from an associated aggregate global financial network (i.e., a single-layer network), long regarded as the standard model for such investigations del Rio-Chanona et al. (2020).
Another study constructed a multiplex network of financial institutions, with layers corresponding to different classes of derivative contracts (e.g., interest rate, credit, foreign exchange), revealing how liquidity shocks propagate across markets Bardoscia et al. (2019). A linked multiplex network incorporating ownership, shared board members, R&D partnerships, and stock correlations among companies in the UK and Germany uncovered a higher level of corporate interconnectedness than previously reported Jeude et al. (2019).
Additional insights have come from analysing the multilayer structure of international trade, where nodes represent countries, layers correspond to industries, and links connect buyers and sellers. A multilayer nestedness metric captured structural organisation more effectively than traditional bipartite models Alves et al. (2019). Similarly, the integration of ownership, innovation collaborations, and board interlocks has shed light on the formation of innovation clusters and their influence on market concentration Glattfelder (2013).
Systemic risk, a core concern in financial stability, is deeply rooted in network interdependencies. Multilayer models have advanced our understanding of how interconnections amplify or mitigate risk Battiston et al. (2012); Caccioli et al. (2014). Studies of global supply chains have also benefited from multilayer perspectives, identifying fragile sectors such as pharmaceuticals, food, and services as critical nodes in system-wide vulnerability Gomez et al. (2020). Lastly, the cryptocurrency ecosystem—comprising thousands of token layers linked by users, exchanges, and dark markets—presents a complex but largely unexplored multilayered network.
Altogether, these examples underscore the promise of multilayer network analysis in uncovering hidden interdependencies, enhancing systemic risk models, and informing economic policy in an increasingly interconnected world.
III.5 Infrastructure networks
Infrastructure systems, such as power grids, transportation networks, and communication systems, are important examples where multilayer networks provide insights into interdependencies and weaknesses. Among these, power grids received significant attention in the early days of the field, particularly after it was shown that interdependencies can make networks more fragile Buldyrev et al. (2010) (see also Section III.1.) This work was motivated by the 2003 blackout in Italy, where failures in power stations led to disruptions in communication networks, which in turn caused further failures in the power grid itself. Subsequent studies confirmed these findings and investigated how various types of interconnections influence robustness Gao et al. (2012).
This research line has since diversified, exploring network fragility from various angles Duan et al. (2019); Artime et al. (2024). For example, some studies have examined cascading failures using load-sharing dynamics Brummitt et al. (2012); Artime and De Domenico (2020), while others have focused on designing inherently more robust networks Reis et al. (2014); Latora and Marchiori (2005). There has also been growing interest in identifying early warning signs of collapse, a challenge relevant not only to infrastructure, but also to biology, economics, and ecology Duan et al. (2019).
More recently, researchers have noted that traditional interdependent network models often assume ”hard failures,” where the collapse of a node in one layer immediately causes failure in a dependent node. Yet, such failures are rare in practice. Instead, attention has shifted toward modelling interdependencies during recovery processes, where one network relies on resources from another to regain functionality Danziger and Barabási (2022).
Transportation systems, much like power grids, have been extensively studied through the lens of multilayer networks Alessandretti et al. (2022). Early work modelled the European air transportation network to assess the impact of flight cancellations on network performance Cardillo et al. (2013). Urban mobility networks further highlight the complexity of multilayer infrastructure in metropolitan areas. Cities rely on multimodal transportation—streets, subways, buses—that together form the backbone of mobility (Fig. 8Aleta et al. (2017); Natera Orozco et al. (2020). Studies have shown that poor synchronisation between modes can lead to significant inefficiencies, including delays and reduced accessibility Gallotti and Barthelemy (2014, 2015). Moreover, navigating these systems can challenge human cognitive limits, making trip planning difficult in highly interconnected urban environments Gallotti et al. (2016).
From a dynamical standpoint, several researchers have explored congestion in multilayer transportation systems Du et al. (2016); Lampo et al. (2021). Notably, increasing speed in one layer—such as subways—can sometimes produce counter-intuitive effects, like congestion at terminal stations due to imbalanced flow dynamics Strano et al. (2015). Expanding the scope beyond transportation modes, urban mobility was analysed by categorizing daily activities (e.g., visiting tourist attractions or sports venues) into distinct layers, offering a multidimensional representation of urban systems Gallotti et al. (2021).
In summary, multilayer network analysis has deepened our understanding of infrastructure systems, from cascading failures in power grids to mobility challenges in urban transport. These insights help identify vulnerabilities, design more resilient systems, and optimise performance, ultimately enhancing the functionality and robustness of essential infrastructure in modern society.
Figure 8: Multilayer representation of Madrid’s public transportation system in 2016. The network is organized into three layers, each representing a different mode of transportation: the top layer (cyan nodes) corresponds to the tram system, the middle layer (orange nodes) to the metro system, and the bottom layer (dark blue nodes) to the bus system. Vertical links connect stops of different modes located within a 150-meter radius, illustrating intermodal connections. Adapted figure with permission from Aleta and Moreno (2019). ©2019 by Annual Reviews.
III.6 Ecological networks
Traditionally, ecological networks have been represented as single-layer networks, focusing on single types of interactions, such as predator-prey or plant-pollinator interactions Tylianakis and Morris (2017). Similarly, although biological communities are intrinsically dynamic, with species and interactions changing over time, temporal dynamics have often been analysed using independent snapshot networks. Thus, ecological systems provide a perfect example of where multilayer networks can play a crucial role by allowing us to simultaneously capture different types of pairwise interactions García-Callejas et al. (2018) or temporal and spatial dimensions, providing a more complete view of ecosystem functioning Hutchinson et al. (2019); Finn et al. (2019).
One study used a five-year dataset on bird-seed dispersal interactions to analyse species turnover and persistence, identifying a stable core of species that consistently supported network structure over time Costa et al. (2020). Similar insights have emerged from studies of plant-pollinator networks that tracked the spatial and temporal persistence of species and interactions Hervías-Parejo et al. (2023), and from work emphasizing the advantages of multilayer approaches over traditional habitat discretisation methods Timóteo et al. (2018).
The common binary classification of species as mutualists or antagonists has also been challenged using multilayer frameworks. In one example, seed viability from faecal samples of wild-trapped mammals showed that some species acted as both antagonists and mutualists for the same plant species, depending on context—revealing a much richer set of interactions than standard models capture Genrich et al. (2017). Multilayer analysis of frugivory and nectarivory interactions has further revealed non-obvious rules of community assembly Mello et al. (2019), while quantifying link strength in networks of pollinators and seed dispersers uncovered that only a subset of individuals truly connects both ecological functions Hervías-Parejo et al. (2020).
Animal social networks also benefit from multilayer analysis, as interactions often span multiple social contexts. A study of baboons with layers representing grooming and association interactions demonstrated that an individual’s centrality ranking can vary dramatically depending on which layer or aggregation is considered Finn et al. (2019). This underlines the importance of selecting relevant interactions when drawing ecological or behavioural inferences.
Recently, a global multilayer modelling framework spanning six ecological functions—pollination, herbivory, seed dispersal, decomposition, nutrient uptake, and fungal pathogenicity—enabled the identification of keystone species, defined by their critical role in maintaining ecosystem function regardless of abundance Hervías-Parejo et al. (2024).
Beyond species interactions, multilayer networks have been applied to physical ecological systems. For instance, river deltas have been modelled as multilayer transport networks Tejedor et al. (2018), and ecosystem restoration has been studied by tracking structural and functional changes in multilayer networks over time Moreno-Mateos et al. (2020). Comparing evolving network states to reference conditions helps assess recovery trajectories.
Despite the promise of multilayer networks in ecology, challenges remain. These include the difficulty of collecting fine-grained interaction data Seibold et al. (2018), the complexity of comparing fundamentally different interaction types Hervías-Parejo et al. (2024), and the ongoing need to develop multilayer metrics tailored for ecological applications Garrido et al. (2023). Addressing these limitations will be key to further advancing ecological insights through multilayer network approaches.
III.7 Science of science
The science of science is the study of the structure and evolution of science via the analysis of data on scientists and their interactions Fortunato et al. (2018); Wang and Barabási (2021). Although the field dates back several decades Price (1963), it has grown rapidly in recent years due to the increasing availability of data on scientific productivity, collaboration, citation, mobility, funding, etc., and to advances in analytical tools, especially in machine learning. Network science has heavily contributed to the development of the field, due to the many networks that can be built from the data.
Multilayer networks have brought added value to this domain by capturing different dimensions of scientific interactions. Early studies showed that combining collaboration networks across journals or merging collaboration and citation networks into a single layer can obscure meaningful correlations Menichetti et al. (2014). A strong correlation between community structures was observed across layers of a multiplex collaboration network of physicists, where each layer represents a different subfield Battiston et al. (2016).
The multilayer framework has also revealed intricate dependencies between collaboration and citation behaviours. For example, papers authored by scientists with a large number of previous collaborators tend to gain citations more quickly, but also fade faster in attention Zingg et al. (2020). Such dependencies can be exploited to develop more accurate generative models of citation networks Nanumyan et al. (2020). Moreover, multilayer networks are better than single-layer ones when it comes to predicting future links, specifically future collaborations between scientists. For this task, besides the collaboration network itself, useful layers may consist of citations between scientists and semantic closeness of the keywords of their papers Tuninetti et al. (2021). Other relevant network types include co-citation and co-venue networks, where two scientists are connected if they cite the same papers or publish in the same journals Pujari and Kanawati (2015).
Ranking scientists by impact is another area enriched by multilayer methods. Traditional indicators such as raw citation counts or the h-index Hirsch (2005) may overlook important structural information. Multilayer networks make it possible to compute composite impact scores that account for multiple dimensions, such as citations and collaborations, either by aggregating layer-specific rankings Pradhan et al. (2017) or through diffusion-based algorithms like multilayer PageRank Omodei et al. (2017). The framework also extends to innovation studies, particularly in analysing patent citation networks. For example, a multiplex network in which layers correspond to different patent-granting offices has been shown to better capture the technological classification of patents than aggregated representations Higham et al. (2022).
Overall, multilayer networks provide a richer, more nuanced view of scientific ecosystems, enhancing our ability to model, predict, and evaluate scientific activity across various dimensions.
III.8 Climate science
Network analysis has become a valuable complement to numerical modelling in climate science Ludescher et al. (2021). In climate networks, nodes typically represent geographic locations, and links indicate similarities between these locations based on relevant physical quantities such as temperature, precipitation, atmospheric pressure, or CO2 concentration. These similarities are quantified using various linear and non-linear measures computed on time series, including Pearson correlation, mutual information, event synchronisation, transfer entropy, partial correlation, or Granger causality.
Most early studies focused on single-variable networks, but recent work has begun to leverage multilayer networks to capture interdependencies among different climate variables or the same variable under different conditions. For instance, two-layer networks constructed from atmospheric pressure fields at different altitudes have revealed significant insights into the general circulation and stratification of the atmosphere Donges et al. (2011).
Multilayer networks also offer powerful tools for assessing climate vulnerability. By integrating climatic and non-climatic factors across multiple layers, researchers can analyse how exposure, sensitivity, and adaptive capacity interact to shape a community’s vulnerability to climate change Debortoli et al. (2018). Similarly, multilayer behavioural networks that integrate supply-side production and demand-side consumption decisions have been used to explore how climate shocks propagate through social and economic systems Naqvi and Monasterolo (2021).
Air pollution is another area where multilayer networks have yielded novel insights. A multilayer network study linking the 500-hPa pressure height (geo-potential height) with surface-level air pollution in China and the United States demonstrated how Rossby waves (large-scale atmospheric waves induced by Earth’s rotation) influence pollution through their effects on cyclone and anticyclone systems Zhang et al. (2019).
The study of CO2 dynamics, a major driver of global warming, has also benefited from multilayer approaches. One interlayer climate network used atmospheric mid-tropospheric CO2 records and ground-level surface air temperature (SAT) data to investigate interactions between CO2 and SAT Ying et al. (2020). This analysis revealed strong interlayer connections, with the most significant links often spanning more than 7,000 km, a distance consistent with the wavelength of atmospheric Rossby waves. A follow-up study removed the influence of Rossby waves to isolate regions where CO2 concentrations are most directly related to SAT variations Ying et al. (2021).
Thus, multilayer network analysis enables a more holistic understanding of the Earth’s climate system by revealing complex interdependencies among climatic variables and between environmental and societal factors. These insights can inform more effective strategies for climate adaptation, risk assessment, and policy development.
III.9 Network medicine
Network medicine applies network science to biological and medical data to better understand, prevent, and treat diseases Barabási et al. (2011). Biological systems encompass a wide variety of networks, including protein–protein interactions (PPIs), metabolic networks, gene regulatory networks, or cell–cell communication networks. Traditionally analysed in isolation, recent years have seen a shift toward integrating diverse omics data (genomics, proteomics, metabolomics, metagenomics, phenomics, and transcriptomics), leveraging multilayer networks to provide a more holistic view of complex biological processes.
Multi-omics integration, especially when coupled with longitudinal data, enables researchers to observe dynamic interactions between biological layers, identify key components of system development, and better understand complex phenotypes. Among biological networks, PPIs are some of the most extensively studied. In multiplex PPI networks, layers may represent interactions in different types of cells (e.g., normal vs. cancerous), different interaction types (e.g., genetic vs. physical), or different data sources (e.g., co-expression, co-annotation)Rai et al. (2017); Yu et al. (2020); Zhao et al. (2016). Such representations have been used to detect biologically enriched modulesMangioni et al. (2018) and to improve the prediction of protein functions.
Figure 9: Multilayer network featuring different omics (genes, proteins, metabolites) and their mutual interactions/associations. Adapted figure with permission from Hammoud and Kramer (2020). ©2020 by Springer Nature.
More frequently, scholars build multilayer networks by combining different omics, as illustrated in Fig.9. In one of the first works, layers consisted of transcription factor co-targeting, microRNA co-targeting, protein-protein interaction, and gene co-expression networks. Communities of genes were extracted by combining partitions found in the individual layers via consensus clustering Lancichinetti and Fortunato (2012). Enrichment analysis on the genes belonging to the same community revealed a set of candidate cancer drivers Cantini et al. (2015b).
Robustness is a prominent feature of most biological systems. By studying the spreading of perturbations in a multilayer network composed of a gene regulatory layer, a protein–protein interaction layer, and a metabolic layer, one can characterise the contribution of each gene to the overall system’s robustness: the most influential genes are enriched in essential and cancer genes Liu et al. (2020). An analysis of the spreading of perturbations through the three layers of a graph representation of the integrative system of E. coli’s gene regulation and metabolism shows that small perturbations, originating in the gene regulatory domain, typically trigger far-reaching system-wide cascades, while small perturbations in the metabolic domain tend to remain more local and trigger much smaller cascades Klosik et al. (2017).
Random walk dynamics offer a powerful framework for analysing multilayer biological systems. For example, random walks across networks of disease-perturbed proteins, drug targets, and biological functions help trace how drug effects propagate Ruiz et al. (2021). Other studies use random walks on multi-omics networks (combining layers from mRNA expression, PPIs, KEGG pathways, and Gene Ontology terms) to map omics molecules to biological functions Bodein et al. (2022). This methodology has also been applied to COVID-19 datasets, enabling insight into disease severity and therapeutic targets Agamah et al. (2024). Single-cell multi-omics multilayer networks have further improved predictions of transcription factor binding sites, regulatory regions, and gene communities Trimbour et al. (2024).
Another area of interest involves disease–disease networks, where nodes are diseases and edges represent functional, semantic, or co-morbidity associations. For instance, two-layer networks combining gene-based and ontology-based similarities have been used to identify diseases with potential shared drug treatments Cheng et al. (2014). Multiplex networks with phenotypic, genetic, pathway-based, and toxicogenomic co-morbidities have revealed the dominant causative mechanisms for hundreds of diseases Klimek et al. (2016a). In another study, a bilayer network of 779 diseases—connected via shared genes and symptoms—enabled the detection of disease communities using multiplex Infomap Halu et al. (2019); De Domenico et al. (2015a). Similarly, during the COVID-19 pandemic, multilayer approaches were instrumental in understanding the disease. The CovMulNet19 network integrated SARS-CoV-2 proteins, host proteins, related diseases, symptoms, and drug targets, and its analysis revealed high similarity between COVID-19 and diseases that are also known patient risk factors Verstraete et al. (2020).
Finally, the growing field of network embeddings has facilitated the integration of heterogeneous biological datasets. By embedding multilayer biological networks into geometric spaces, researchers have improved tasks like drug–target interaction prediction and biomarker identification Zitnik and Leskovec (2017); Jagtap et al. (2022) (see Section II.3.3).
In short, multilayer networks are proving essential for capturing the complexity of biological systems and improving our understanding of disease mechanisms, interactions, and potential treatments.
III.10 Network neuroscience
Neuroscience is fundamentally centred on the study of the brain as a vast, interconnected system of neurons and brain regions. This field encompasses not only the structural and functional organisation of the brain but also the dynamic processes that underpin cognition, behaviour, and neurological health. Networks provide a natural framework for capturing these complexities by representing the brain as a system of nodes and edges. Over the past decade, network neuroscience Bassett and Sporns (2017); Barabási et al. (2023) has emerged as one of the most exciting and influential fields, offering new insights into the brain’s organisation and its intricate functions.
Brain networks Bullmore and Sporns (2009) are mathematically modelled systems where nodes represent either neurons or specific brain regions, while edges capture various forms of connectivity. These connections may represent structural connectivity, which reflects the physical pathways linking brain regions and is often derived from diffusion MRI or diffusion tensor imaging. Alternatively, they may capture functional connectivity, which describes statistical relationships in neural activity between regions, as observed through functional MRI, electroencephalography, or magnetoencephalography. A third type, effective connectivity, goes a step further by modelling causal interactions and directional relationships between regions, often using techniques like Granger causality or dynamic causal modelling Nicosia et al. (2013b). Together, these distinct forms of connectivity provide a rich and multifaceted view of the brain’s organisation and activity.
Building on this foundation, multilayer networks offer a sophisticated extension for exploring brain connectivity and dynamics De Domenico (2017); Presigny and De Vico Fallani (2022). These models expand the traditional network framework by incorporating multiple interconnected layers of nodes and edges, each representing a different dynamic or scale of analysis Betzel and Bassett (2017). Within an individual, for example, these layers can correspond to distinct types of connectivity, such as structural and functional networks, with interlayer links capturing relationships between them Vaiana and Muldoon (2020). Similarly, temporal transitions in dynamic functional connectivity can be represented as changes within and across layers or states Sporns et al. (2021). Beyond individuals, multilayer networks also enable the comparison of connectivity patterns across populations by modelling interlayer relationships that represent similarities between individuals or groups Wilson et al. (2020); Simas et al. (2015). This versatility allows multilayer networks to bridge gaps between data types and scales, revealing insights that would be inaccessible using single-layer approaches Battiston et al. (2018); Mandke et al. (2018); Presigny et al. (2024).
The ability to integrate and analyse these diverse sources of data has led to important studies in areas such as human learning Bassett and Mattar (2017), individual differences Betzel et al. (2019), the diagnosis and understanding of neurological and mental disorders Stam (2014), among others like the relation between social environment and variation in brain network organisation Merritt et al. (2024).
Additionally, even though traditional network models focus mainly on pairwise relationships, higher-order interactions can be pulled into the multilayer network paradigm. For instance, one could consider multiple layers describing different orders of interaction, an approach that has contributed to new advances in the understanding of human brain evolution and cognition Luppi et al. (2022), and the relation between brain activity and behaviour Santoro et al. (2024). In addition, higher-order structures can help uncover the relationship between structure and function in the brain Battiston et al. (2017).
In summary, the multilayer networks framework represents a frontier in network neuroscience, enabling researchers to integrate diverse data types and explore the brain’s multidimensional nature. Continued methodological advances will undoubtedly unlock clearer insights into brain function and dysfunction, fostering progress in both neuroscience and clinical applications.
IV Data & Tools
Given the growing interest in network science, a wide array of resources has become available. Yet, despite substantial progress, challenges remain in terms of data availability, standardisation, and tool interoperability. This section highlights the most relevant public data repositories and software libraries that support advanced network analysis and visualisation.
IV.1 Data
Network data comes from a wide range of disciplines such as biology, sociology, economics, neuroscience, and more. While numerous repositories exist for general-purpose network datasets, multilayer and temporal network datasets are comparatively less common and often tailored to specific domains. Furthermore, the lack of standardised formats can hinder reproducibility and comparative studies across methods.
Nevertheless, various curated and open-access repositories offer high-quality datasets that have become benchmarks in the field. Table 1 lists key sources of network datasets across domains. It is worth noting that the table only includes major and widely used repositories. Many individual researchers also make their datasets publicly available on personal or project websites. While this is an invaluable contribution to the community, it also highlights the fragmentation of the field and the lack of consistent standards for sharing and documenting network data, although there is a lot of effort to improve this situation.
Repository Networks Reference
Netzschleuder 163735 Peixoto (2020)
Network Repository 5664 Rossi and Ahmed (2015)
KONECT 1326 Kunegis (2013)
ICON 704 174
SNAP 127 Leskovec and Krevl (2014)
UCI 35 DuBois (2008)
Table 1: Examples of publicly available network datasets, sorted by number of networks available as of September 2025. Some repositories include multilayer, temporal, and higher-order networks.
While many general-purpose datasets are readily accessible and well-documented, multilayer and higher-order data remain underrepresented. A key issue is the lack of standardised file formats to represent these complex structures, which can complicate interoperability between tools and comparisons across studies.
IV.2 Software Tools
The development of dedicated software libraries has significantly accelerated the field of network science. A wide variety of tools exist for analysis, simulation, and visualisation of networks. Table 2 summarizes major software libraries, including their main focus and programming language. It is important to note that this overview focuses on general-purpose tools for the structural analysis and visualisation of complex networks. Tools primarily focused on specific functional processes or computational workflows, such as those for graph neural networks PyTorch Foundation (2025) or epidemic simulations Czuba et al. (2024); Gozzi et al. (2025); Miller and Ting (2019), are not included.
However, no single tool dominates the field. Even for general-purpose networks, researchers choose between different libraries based on project requirements. For example, while networkX is widely used for its ease of use and versatility, alternatives such as igraph or graph-tool offer superior performance on large networks due to optimised C/C++ backends. Thus, selecting the appropriate tool often involves trade-offs between ease of use, scalability, feature set, and visualisation capabilities.
Type Name Main Focus Language Reference
Analysis DyNetX Dynamic networks Python Rossetti (2020)
HAT Higher-order networks Python, MATLAB Pickard et al. (2023)
graph-tool General-purpose networks Python Peixoto (2014)
HGX Higher-order networks Python Lotito et al. (2023)
HyperNetX Higher-order networks Python Praggastis et al. (2024)
iGraph General-purpose networks Python, R, C Csárdi and Nepusz (2006)
MultilayerGraphs.jl Multilayer networks Julia Moroni and Monticone (2022)
multinet Multilayer networks R Magnani et al. (2021b)
NetworKit Large-scale networks Python Research Group Modeling and Analysis of Complex Systems (2025)
networkX General-purpose networks Python Hagberg et al. (2008)
NDlib General-purpose networks Python Rossetti et al. (2018)
pathpy Higher-order temporal networks Python Hackl et al. (2021)
pymnet Multilayer networks Python Nurmi et al. (2024)
pyunicorn General-purpose networks Python Donges et al. (2015)
Reticula Higher-order temporal networks Python Badie-Modiri and Kivelä (2023)
SimpleHypergraphs.jl Higher-order networks Julia Spagnuolo et al. (2020)
straph Temporal networks Python Latapy et al. (2018)
tacoma Temporal networks Julia Maier and Schlosser
TGLib Temporal networks Python, C++ Oettershagen and Mutzel (2022)
tnetwork Temporal networks Python Remy and Jorquera
XGI Higher-order networks Python Landry et al. (2023)
Visualisation Cytoscape General-purpose networks GUI Shannon et al. (2003)
Gephi General-purpose networks GUI Bastian et al. (2009)
ggraph General-purpose networks R Pedersen (2024)
Graphia General-purpose networks GUI Freeman et al. (2022)
MultiNet Multivariate networks Web-based Wilburn and Lex (2023)
multiNetX Multilayer networks Python Kouvaris
muxViz Multilayer networks R, Web-based De Domenico (2022)
Netminer General-purpose networks GUI Ghim et al. (2014)
Netwulf General-purpose networks Python, Web-based Aslak and Maier (2019)
NodeXL General-purpose networks Excel 251
Pajek General-purpose networks GUI Batagelj and Mrvar (2004)
py3Plex Multilayer networks Python Skrlj et al. (2019)
Sigma.js General-purpose networks JavaScript 302
Tulip General-purpose networks GUI Auber et al. (2012)
Table 2: Libraries and tools for network science analysis and visualisation, with support for multilayer, temporal, and higher-order networks. “Main focus” refers to the area where the software is primarily applied; however, some tools can support multiple network types.
V Outlook
The past decade has shown that multilayer networks are not merely a convenient generalization of graph theory but a fundamentally new modelling language. The field has now reached a stage where the central question is no longer how to adapt single-layer tools, but how to identify the truly multilayer mechanisms that cannot be reduced away. Several theoretical directions emerge from this perspective. One concerns the search for minimal descriptions: determining when layers encode irreducible information and when they can be compressed. Another direction concerns the interplay between multilayer topology and cross-layer dynamical coupling, where instabilities, hybrid phase transitions, metastability, or coexistence of dynamical regimes can arise due to the multiplex nature of the system. Understanding the general principles that govern these phenomena remains a major open challenge. At the same time, the field’s inherently multidisciplinary nature complicates the picture: multilayer approaches are used in physics, biology, economics, computer science, and the social sciences, often with different assumptions, terminologies, and modelling priorities. This diversity makes communication across communities difficult, yet it also underscores the value of multilayer networks as a common representational language capable of unifying disparate viewpoints.
At the methodological frontier, several additional opportunities demand attention. The first involves integrating multilayer representations with temporal and higher-order structures. Many real systems combine heterogeneous interaction types with time-varying group interactions, but current modelling frameworks still tend to treat these features separately. Developing unified theoretical tools capable of capturing polyadic interactions, memory effects, and cross-layer dependencies simultaneously is an important next step. A second direction concerns inference, where the rise of generative models, statistical reconstruction techniques, and machine-learning-based embeddings is beginning to provide new ways to infer hidden layers and mesoscale structures. Yet, understanding the limits of inference remains unresolved. A third challenge lies in controllability and optimisation: while single-layer structural controllability is now well understood, its multilayer analogue is largely unexplored, especially when layers interact through non-linear or time-dependent couplings. Finally, the rapidly growing integration between network science and artificial intelligence opens a promising but underdeveloped path. Deep learning architectures are themselves multilayered systems whose internal dynamics could benefit from tools developed in multilayer network science, potentially contributing to interpretability and robustness in AI.
Across domains, applications of multilayer network science are entering a phase in which data and theoretical tools are finally becoming aligned. A key opportunity lies in modelling coupled processes in the wild: infrastructure failures intertwined with cyber systems, economic–epidemic feedbacks, multi-omics interactions, or the joint evolution of online and offline social behaviour. These problems require models that simultaneously encode structural heterogeneity, temporal evolution, and cross-layer feedback, which are precisely the strengths of the multilayer paradigm. A central challenge, however, is that multilayer data and tools originate from many different fields, each with its own conventions, formats, and epistemic goals. This heterogeneity makes it difficult to compare studies or integrate datasets, yet it also reflects the breadth of the systems that multilayer network science can represent. As common standards and interoperable software mature, they provide a platform for genuine cross-disciplinary exchange, helping the field not only accumulate data but also communicate across communities. In scientific domains ranging from ecology to neuroscience, such integrative approaches are beginning to reveal how processes at different scales coordinate, interfere, or reinforce each other.
Looking ahead, progress will depend critically on the availability of high-resolution multilayer data, standardised formats, and interoperable tools, but also on conceptual clarity about what it means for a system to be truly multilayer. If these developments continue in parallel, multilayer network science is poised to move from descriptive analyses to predictive, actionable models capable of guiding interventions, optimising coupled infrastructures, and informing policy in domains where interdependence is no longer optional to ignore. As the complexity of modern systems continues to grow, the multilayer perspective will remain essential for understanding the intertwined processes that characterise the natural and engineered world.
Acknowledgements.
All authors acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (Award #1927425 and #1927418). We thank M. Clarin from COSNET Lab-BIFI and N. Samay from NetSI for helping with the figures. A.S.T. acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UID/408/2025. A.A. acknowledges support from the grant RYC2021‐033226‐I funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. A.A. and Y.M. were partially supported by the Government of Aragon, Spain, and ERDF ”A way of making Europe” through grant E36-23R (FENOL), and by Grant No. PID2023-149409NB-I00 from Ministerio de Ciencia, Innovación y Universidades, Agencia Española de Investigación (MICIU/AEI/10.13039/501100011033) and ERDF ”A way of making Europe”. G.F.A was partially supported by the São Paulo Research Foundation (FAPESP), through grants 2024/16711-8 and 2025/04409-8. A.D.-G. and O.A. acknowledge support from the Spanish Grant No. PID2021-128005NB-C22, funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe;” and from Generalitat de Catalunya (No. 2021SGR00856). M.S. acknowledges support from Grants No. RYC2022-037932-I and CNS2023-144156 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. M.K. acknowledges support from the ANR funded DATAREDUX project (ANR-19-CE46-0008); the National Laboratory for Health Security, Alfréd Rényi Institute of Mathematics, RRF-2.3.1-21-2022-00006; and the WWTF funded MOMA project (10.47379/ESS22032). J.K. gratefully acknowledges partial support by ERC grant No. 810115-DYNASNET. G.P acknowledges support from the European Research Council (ERC) Consolidator Grant under the European Union’s Horizon Europe programme (grant agreement No. 101171380, project RUNES) and the MSCA Doctoral Network BeyondTheEdge(Grant no. 101120085).